# MOOSE-Chem: Large Language Models for Rediscovering Unseen Chemistry Scientific Hypotheses

### v1

@@ 1 Introduction
We design experiments with the benchmark to test the three fundamental questions, and find that LLMs are highly capable. We also test MOOSE-Chem with the benchmark, mimicking the setting to run it in the wild by only giving the background and a corpus of up to 3000 chemistry papers to select inspiration. Even in this challenging setting, MOOSE-Chem can still rediscover many hypotheses with very high similarity with the ground truth ones, covering the main innovations.
@@ 2 Related Work
Zhong et al. (2023) work on finding the difference between two corpora to propose hypotheses, but their evaluation is based on selecting hypotheses that do not require expert knowledge for Turkers to judge, thus cannot lead to a novel discovery. Wang et al. (2024) try to utilize LLMs to discover novel NLP hypotheses, and find the hypotheses still fall far behind scientific papers in terms of novelty, depth, and utility. Yang et al. (2024b) show that LLMs can generate novel and valid enough hypotheses evaluated by PhD students, but they only focus on social science. FunSearch (Romera-Paredes et al., 2024) can discover specific solutions for mathematical conjecture but can’t discover new math theorems. Qi et al. (2024) analyzes LLM’s ability for scientific discovery in the biomedical domain by directly generating hypotheses with only the research background. Boiko et al. (2023); Baek et al. (2024); Li et al. (2024); Lu et al. (2024) focus on subsequent steps for scientific discovery, mainly developing and conducting experiments. Sprueill et al. (2023; 2024) focus on catalyst discovery, but their evaluation relies on whether can rediscover existing commercially used catalysts, which might cause data contamination problem. Tshitoyan et al. (2019) show that word embedding obtained from large-scale chemistry literature can recommend materials for functional applications years before their discovery by controlling the date of the training corpus. (Xie et al., 2024) predict emerging thermoelectric materials by summarizing the sentiment in the existing literature.
@@ 3 Benchmark Construction
Table 2 and Table 2 shows the statistics of the benchmark in terms of chemistry category and publication venue. Many collected benchmark papers are related to material science (around 13 papers). Material science is a sub-category of chemistry and can belong to the categories in Table 2, such as polymer material and organic material.
@@ 4.1 Fundamental Assumption and Following Decomposition
In other words, given existing knowledge in the background, a majority of chemistry research is about searching knowledge that previously not known to be related to the background but in fact can assist the background, then associate the searched knowledge and the background knowledge in a reasonable way to compose a hypothesis. For example, the proposal of backpropagation can be seen as a hypothesis. In this case, the background knowledge is multi-layer logistic regression, and the searched knowledge is the “chain-rule” in mathematics.
@@ 4.2.1 The General Picture
Our answer is, at least we might need to try multiple times and various ways to leverage the chain rule for multi-layer logistic regression. With this motivation, we develop a simple Evolutionary Algorithm (EA) based method, shown in Figure 1.
@@ 5 Investigation on Fundamental Questions
Here we investigate the question (denoted as $Q1$): “whether LLM can identify inspiration papers which is unknown to be able to associate with the background (or at least unknown to associate in a certain way) but in fact have the potential to help with the background question?”.
@@
Table 3 shows the main experiment results for $Q1$. The Hit Ratio is surprisingly high. Particularly, more than 75% of ground truth inspirations are selected even only to select 4% papers from the chemistry literature corpus. It seems that LLMs are quite capable of finding inspiration papers that are unknown to be able to associate with the background but, in fact, have the potential to help with the background question. It means our bold assumption in § 4.2.2 that “the most advanced LLMs might already know lots of knowledge pairs that are able to associate to create novel knowledge, where the knowledge pairs are not known by any scientist to be related” is possible to be true.
@@
Table 7 shows the results. For each $b$, the top two $h$ with the highest MS by GPT-4o are selected for expert evaluation (by two chemistry PhD students). It indicates that LLMs are quite capable to associate known knowledge into an unknown knowledge that has high probability to be valid (very close to $oh$). In addition, providing a survey can assist the new knowledge discovering process.
@@ 5.3 How Well can LLMs perform $R(h)$?
To investigate $Q3$, we run MOOSE-Chem with every $b$ from the benchmark; $|I|=300$, containing all the ground truth $i$. Every $h$ is given a rating $r=R(h)$, and is ranked based on $r$. For every generated $h$, we get the number of ground truth $i$ it leveraged (#Matched $i$), and assign it with a GPT-4o evaluated MS (here MS is -1 means this $h$ has not used any ground truth $i$).
@@
Table 9 shows the relation between the GPT-4o evaluated MS and the average ranking ratio. For $h$ with a positive MS, there is a trend that the higher the MS, the better the average rank ratio (if MS $\in$ [2,4]). However, the disadvantage of those $h$ without a positive MS is not very significant. It seems that LLMs have a certain ability to rank good $h$ higher. But it is not very significant. A part of the reason might be that those $h$ generated without groundtruth $i$ are also in high quality.
@@ 6.1 Baselines
is a hypothesis discovery framework for the social science domain. It leverages LLMs to retrieve inspirations and uses self-refine (Madaan et al., 2023) to improve the validness, novelty, and clarity aspects. The difference is that (1) it does not adopt the mutation and recombination step to better associate background and inspiration; (2) it only retrieves one step of inspiration.
@@
is a hypothesis discovery framework for the NLP domain. It relies on semantic and citation neighbors to retrieve information to assist the background. As a result, the retrieved information could be very related to the background that might not be able to serve as an inspiration. To make the generated hypothesis more novel, it adopts self-refine to focus on improving the novelty aspect of the generated hypothesis. Here we implement SciMON with LLM-based inspiration retrieval, the same as MOOSE-Chem. Table 3 shows that the recall rate of LLM-based retrieval is 83.7%.
@@
work on hypothesis discovery in Biomedical domain. It retrieves information pertinent to the keywords in the background to generate hypotheses. As a result, the retrieved information might compose of a background survey, but not as inspiration. Self-refine is also adopted.
@@ 6.2 Results
Table 10 shows the baseline results and the ablation study of MOOSE-Chem. It indicates that both mutation & recombination and the multi-step designs can significantly improve the best-performing $h$. Mutation & recombination leads to a drop of Average MS; we attribute the reason to that the mutation step forces LLMs to generate $h$ different from previous $h$ mutations from the same $b$ and $i$, and therefore might generate many $h$ that do not make lots of sense. The assigned MS to these mutation $h$ is low, and therefore lower down the Average MS.
To better understand the performance of MOOSE-Chem in this real copilot setting, for each $b$ we select 4 generated $h$ with the highest MS by GPT-4o. Two experts evaluate these $h$ again in terms of MS. Table 11 shows the expert evaluation results. Here the top MS is the highest MS for each $b$, out of the 4 expert evaluated $h$ for this $b$. Note that MS rated as three is already very high. Illustrated in Table 6, it means the generated $h$ by MOOSE-Chem (that has not seen $h$) in the real copilot setting covers two main innovations of the hypothesis published in Nature, Science or a similar level.
@@ 8 Conclusion
We investigate this central question: ‘Can LLMs automatically discover novel and valid chemistry research hypotheses (even in the Nature level) given only a chemistry research background (consisting of a research question and/or a background survey), without limitation on the domain of the research question?”. We propose a fundamental assumption to break up the central question into three smaller ones and investigate LLM’s ability on each of them. To do it, we construct a benchmark consisting of chemistry papers published and only be public in 2024. We also develop a multi-agent framework consisting of three stages reflecting the three smaller questions. Experiments show that the framework (runs in a copilot in-the-wild setting) can rediscover many hypotheses with very high similarity with the ground truth ones, covering the main innovations.

### v2

@@ 1 Introduction
We design experiments with the benchmark to test the three fundamental questions, and find that LLMs are highly capable. We also test MOOSE-Chem with the benchmark, mimicking the setting to run it in the wild by only giving a background and a corpus of up to 3000 chemistry papers to select inspiration. Even in this challenging setting, MOOSE-Chem can still rediscover many hypotheses with very high similarity with the ground truth ones, covering the main innovations.
@@ 2 Related Work
Zhong et al. (2023) work on finding the difference between two corpora to propose hypotheses, but their evaluation is conducted by Turkers, which cannot lead to a novel discovery. Wang et al. (2024) try to utilize LLMs to discover novel NLP and biochemical hypotheses, and find the hypotheses still fall far behind scientific papers in terms of novelty, depth, and utility. Yang et al. (2024b) first show that LLMs can generate novel and valid enough hypotheses evaluated by PhD students, but they only focus on social science. FunSearch (Romera-Paredes et al., 2024) can discover specific solutions for mathematical conjecture but can’t discover new math theorems. Qi et al. (2024) analyzes LLM’s ability for scientific discovery in the biomedical domain by directly generating hypotheses with only the research background. Boiko et al. (2023); Baek et al. (2024); Li et al. (2024); Lu et al. (2024) focus on subsequent steps for scientific discovery, mainly developing and conducting experiments. Sprueill et al. (2023; 2024) focus on catalyst discovery, but their evaluation relies on whether can rediscover existing commercially used catalysts, which might cause data contamination problem. Kumar et al. (2024) compare different LLMs on scientific discovery on different disciplines. Tshitoyan et al. (2019) show that word embedding obtained from large-scale chemistry literature can recommend materials for functional applications years before their discovery by controlling the date of the training corpus. (Xie et al., 2024) predict emerging thermoelectric materials by summarizing the sentiment in the existing literature.
@@ 3 Benchmark Construction
Table 2 and Table 2 show the statistics of the benchmark in terms of chemistry category and publication venue. Material science is a sub-category of chemistry and can belong to the categories in Table 2, such as polymer material and organic material. Around 13 collected benchmark papers are inside the material science domain. Beyond them, more papers have intersections with material science. In this paper, we target both chemistry and material science, but for simplicity, we only refer to them as chemistry in this paper.
@@ 4.1 Fundamental Assumption and Following Decomposition
In other words, given existing knowledge in the background, a majority of chemistry research is about searching knowledge that previously not known to be related to the background but in fact can assist the background, then associate the background knowledge and the searched knowledge in a reasonable way to compose a hypothesis. For example, the proposal of backpropagation can be seen as a hypothesis. In this case, the background knowledge is multi-layer logistic regression, and the searched knowledge is the “chain-rule” in mathematics.
@@ 4.2.1 The General Picture
Our answer is, at least we might need to try multiple times and various ways to leverage the chain rule for multi-layer logistic regression. With this motivation, we develop a simple Evolutionary Algorithm (EA) based method, shown in the top-right of Figure 1.
@@ 5 Investigation on Fundamental Questions
Here we investigate the question (denoted as $Q1$): “whether LLM can identify inspiration papers which are unknown to be able to associate with the background (or at least unknown to associate in a certain way) but in fact can associate with the background to create novel knowledge?”.
@@
Table 3 shows the main experiment results for $Q1$. The Hit Ratio is surprisingly high. Particularly, more than 75% of ground truth inspirations are selected even only to select 4% papers from the chemistry literature corpus. It seems that LLMs are quite capable of finding inspiration papers that are unknown to be able to associate with the background but in fact can associate with the background to create novel knowledge. It means our bold assumption in § 4.2.2 that “the most advanced LLMs might already know lots of knowledge pairs that are able to associate to create novel knowledge, where the knowledge pairs are not known by any scientist to be related” is possible to be true.
@@
Table 7 shows the results. For each $b$, the top two $h$ with the highest MS by GPT-4o are selected for expert evaluation (by two chemistry PhD students). It indicates that LLMs are quite capable to associate known knowledge into an unknown knowledge that has high probability to be valid (very close to $oh$). In addition, providing a survey can assist the new knowledge discovering process. We discuss the agreement between GPT-4o based evaluation and expert evaluation in § A.7.
@@ 5.3 How Well can LLMs perform $R(h)$?
To investigate $Q3$, we run MOOSE-Chem with every $b$ from the benchmark; $|I|=300$, containing all the ground truth $i$. Every $h$ is given a rating $r=R(h)$, and is ranked based on $r$. For every generated $h$, we get the number of ground truth $i$ it leveraged (#Matched $i$), and evaluate it with a GPT-4o evaluated MS (here MS is -1 means this $h$ has not used any ground truth $i$).
@@
Table 9 shows the relation between the GPT-4o evaluated MS and the average ranking ratio. For $h$ with a positive MS, there is a trend that the higher the MS, the better the average rank ratio (if MS $\in$ [2,4]). However, the disadvantage of those $h$ without a positive MS is not very significant. It seems that LLMs have a certain ability to rank good $h$ higher. But it is not sure how significant it is, because a part of the reason of this results is that those $h$ generated without groundtruth $i$ could be also in high quality.
@@ 6.1 Baselines
is a hypothesis discovery framework for the general social science domain. It leverages LLMs to retrieve inspirations and uses self-refine (Madaan et al., 2023) to improve the validness, novelty, and clarity aspects. The difference is that (1) it does not adopt the mutation and recombination step to better associate background and inspiration; (2) it only retrieves one step of inspiration.
@@
is a hypothesis discovery framework for the NLP and biochemical domain. It relies on semantic and citation neighbors to retrieve information to assist the background. As a result, the retrieved information could be very related to the background that might not be able to serve as an inspiration. To make the generated hypothesis more novel, it adopts self-refine to focus on improving the novelty aspect of the generated hypothesis. Here we implement SciMON with LLM-based inspiration retrieval, the same as MOOSE-Chem. Table 3 shows that the recall rate of LLM-based retrieval is 83.7%.
@@
work on hypothesis discovery in the biomedical domain. It retrieves information pertinent to the keywords in the background to generate hypotheses. As a result, the retrieved information might compose of a background survey, but not as inspiration. Self-refine is also adopted.
@@ 6.2 Results
Table 10 shows the baseline results and the ablation study of MOOSE-Chem. It indicates that both mutation & recombination and the multi-step designs can significantly improve the best-performing $h$. Mutation & recombination leads to a drop of Average MS compared to the MOOSE baseline; we attribute the reason to that the mutation step forces LLMs to generate $h$ different from previous $h$ mutations from the same $b$ and $i$, and therefore might generate many $h$ that do not make a lot of sense. The assigned MS to these mutation $h$ is low, and therefore lower down the Average MS.
To better understand the performance of MOOSE-Chem in this real copilot setting, for each $b$ the top 4 generated $h$ with the highest MS by GPT-4o are evaluated again by two experts in terms of MS. Table 11 shows the expert evaluation results. Here the top MS is the highest MS for each $b$, out of the 4 expert evaluated $h$ for this $b$. Note that MS rated as three is already very high. Illustrated in Table 6, it means the generated $h$ by MOOSE-Chem (that has not seen $h$) in the real copilot setting covers two main innovations of the chemistry hypothesis, which is published in Nature, Science or a similar level.
@@ 8 Conclusion
We investigate this central question: “Can LLMs automatically discover novel and valid chemistry (including material science) research hypotheses (even those which deserve a publication in Nature, Science, or a similar level) given only a chemistry research background (consisting of a research question and/or a background survey), without limitation on the domain of the research question?”. We propose a fundamental assumption to break up this seemingly-impossible-to-solve central question into three smaller, more practical and executable fundamental questions. Then, we investigate LLM’s ability on each of them. To do it, we construct a benchmark consisting of chemistry and material science papers published and only be public in 2024. We also develop an LLM-based multi-agent framework consisting of three stages reflecting the three smaller fundamental questions. Experiments show that the framework (runs in a copilot in-the-wild setting, with LLMs with training data up to October 2023) can rediscover many hypotheses with very high similarity with the ground truth ones, covering the main innovations.